OpenStudy (anonymous):

hey @lol23 Solve. m – 4 = –8 A. –12 B. –4 C. 4 D. 12

Nnesha (nnesha):

4 is minus on left side you have to move -4 to the right so add 4 both side

OpenStudy (anonymous):

so b

Nnesha (nnesha):

yes that's right

OpenStudy (anonymous):

I gotta go to sleep sorry its 2:00 pm

OpenStudy (phebe):

of the equation : m-4-(-8)=0 Step by step solution : Step 1 : Equation at the end of step 1 : Step 2 : Solve m+4 = 0 Solving a Single Variable Equation : 2.1 Solve : m+4 = 0 Subtract 4 from both sides of the equation : m = -4 ANSWER IS: m = -4

OpenStudy (phebe):

So B is right

OpenStudy (anonymous):

where do u live@lol23 g'night

Nnesha (nnesha):

hmm keep it simple :)

OpenStudy (danjs):

isn't 2 am an hour east of eastern time zone, somewhere in the atlantic, or nova scotia or something